OpenStudy (anonymous):

3^x=4 =7^x-1 . .. how do i solve these kind of problems?

OpenStudy (anonymous):

it looks more like this \[3^{x+4}=7^{x-1}\]

OpenStudy (anonymous):

please help :(

OpenStudy (campbell_st):

ok... use logarithms so taking the log of both sides... doesn't matter if its base e or base 10 \[\log(3^{x +1} )= \log(7^{x -1})\] which can be simplifed using log laws to \[(x + 1)\log(3) = (x - 1)\log(7)\] distributing and collecting like terms you will get \[Log(3) + \log(7) = x \log(7) - x \log(3)\] factor the right hand side \[\log(3) + \log(7) = x( \log(7) - \log(3))\] and divide both sides \[x = \frac{\log(3) + \log(7)}{\log(7)- \log(3)}\] it can be simplified by using log laws for multiplication and addition to \[x = \frac{\log(21)}{\log(\frac{7}{3} )}\] it just depends if you need an exact value answer or an answer to several decimal places.. Hope this helps